OpenStudy (planxavighost):

Which equation evaluates the following problem? Sara saved $100.00 to buy a cell phone and start a monthly plan. The bill for the first month includes everything. She was charged for the cell phone which cost $50.00, the monthly fee of $35.00, and $2.00 for each additional minute used. Sara spent 7 additional minutes that first month talking on the phone. Will she have enough money to pay her phone bill?

OpenStudy (acespeedfighter):

idk much about this sorry :?

OpenStudy (planxavighost):

thanks for trying

OpenStudy (zpupster):

write equation 50.00 and we will add 35.00 for monthly fee now we have 7 add. min. multiplied by 2.00 so 50.00 + 35.00 + (2*7) = ?? is this no. greater than the 100.00 she saved

OpenStudy (planxavighost):

so the answer is?

OpenStudy (planxavighost):

@zpupster

OpenStudy (zpupster):

do the math. we will see if your answer is correct

OpenStudy (mysticforce21):

no she won't

OpenStudy (planxavighost):

@Mysticforce21 what is the answer?

OpenStudy (akshat):

Answer is she would be able as she has already spend $85(55+25) and if she talks 7 min more then 7*2 = 14 so total money spend +85+14= 99 ans he saved 100 so she has 1 $ left with her

OpenStudy (planxavighost):

the answer choices are A.182 B.215 C.253 D.398

OpenStudy (planxavighost):

Solve. Be sure to use the order of operations. 54 + 28 × 5 - 12 =

OpenStudy (mysticforce21):

the answer is sarah will not have enough money to pay her cell phone bill

OpenStudy (mysticforce21):

The answer is 182 ok